OpenStudy (anonymous):

find the solution of this system of equations spate the x- and y- values with a comma x=-10+y -6x+8y=22

OpenStudy (anonymous):

do you know how to do this problem?

OpenStudy (anonymous):

sorry let me rephrase do you know how to start the problem?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok so first you have x = -10 + y and -6x + 8y = 22 right? you need to isolate the variables in one of the equations. it has already been done for you. do you know which one it is?

OpenStudy (anonymous):

ive never taken this class before in highschool and im doing it now for summer school im going to be a senior this year and I don't understand it really math is my weakness

OpenStudy (anonymous):

not sure how to do it

OpenStudy (anonymous):

thats ok! im here to help you! :) so what it means to isolate a variable is that a variable is a letter in any given equation. got this so far?

OpenStudy (anonymous):

i kinda do

OpenStudy (anonymous):

so any letter in the alphabet can be a variable in an equation. for example: 1 + 2 = p 3 + 4 = g 5 + 6 = a so the p, g, and a are the variables in these 3 equations

OpenStudy (anonymous):

but that's not the question i asked

OpenStudy (anonymous):

yes i know that is to help ease you into the problem because you said you only sorta understood so i tried to make it a little bit easier to understand

OpenStudy (anonymous):

so do you understand about variables now?

OpenStudy (anonymous):

seriously?

OpenStudy (anonymous):

you know what? you do it ok? i was just trying to teach you. you asked for someone to teach you and that's what i was trying to do. you think you are all inferior and everything because you are going to be in senior of high school. lemme break it to ya kid... high schoolers don't do this kind of math... 6th graders do. i bet you u can ask any sixth grader if they can do this and they would be able to do this in a minute or less. good luck in life lauren cuz u sure would need it.

OpenStudy (anonymous):

excuse you im not a kid ok im 18 and this is 11th grade math geez kiddo